OpenStudy (anonymous):

One pipe can fill a pool 1.25 times faster than a second pipe. When both pipes are opened, they fill the pool in five hrs. how long would it take to fill the pool in only the slower pipe is used.

OpenStudy (anonymous):

I just need a concept(formula) no need for the answer.. because I should be the one should find it. thanks.

OpenStudy (anonymous):

x hours to fill the pool for fast pipe, 1.25x hours for the slower pipe Together: first pipe in 1 hour + second pipe in 1 hour = 1/5 of the pool 1/x + 1/(1.25x)= 1/5...solve for x, to find the slower pipe take 1.25 time x

OpenStudy (anonymous):

thanks a lot.